OpenStudy (anonymous):

Hopelessly stuck on a problem, any help would be appreciated! Problem - Produce a graph of the line given the equation: 1/4x - 3/5y = 12 Here's what I've worked out so far: - 3/5y = 12 - 1/4x y = 12 - 1/4x * - 5/3 y = 5x/12 - 20 I don't know if i've made a mistake to this point but from here I've tried using a table to find the points to graph but I'm not coming up with anything that makes sense.

OpenStudy (anonymous):

You can multiply all the terms in the equation by any number and the equation will still remain true. In this case, you'll probably want to multiply by 20 to get rid of the fractions. 20 * (1/4 x - 3/5 y = 12) results in 20/4 x - 60/5 y = 240 or 5x - 12y = 240 That might make it a little easier. If you need further help, just ask.

OpenStudy (anonymous):

The y-intercept will be where x=0. So, in your equation y=5x/12 - 20, making x=0 results in y = -20. You can plot that point (at 0, -20).

OpenStudy (anonymous):

So would that make the points (0, -20) (12, -15) and (24, -10)?

OpenStudy (anonymous):

It sure would!

OpenStudy (anonymous):

lol, thanks for the help!